# An effective gauge field theory of the nucleon interactions

### v2

@@ 1 Introduction
One of the first theories of nucleon-nucleon interactions was the Yukawa theory [1]. Subsequently, the meson exchange potentials between nucleons began to be intensively constructed and studied. Various variants of potentials were obtained, which made it possible to understand many properties of nuclei, to quantitatively describe data on cross sections, phases, and scattering lengths. A number of specific potentials have received much attention, such as the Moscow potential [2], the Argonne potential [3] or later ones built on the basis of field theories, such as the Bonn potential [4], the Nijmegen potential [5, 6] and others. Review article [7] provides a detailed overview of the history and current state of research on internucleon potentials.
The idea of exchange interaction, the exchange of pi-mesons and other mesons with masses less than the mass of the nucleon, turned out to be exceptionally fruitful not only for describing nuclear forces. Yukawa-type interactions formed the basis for describing the interaction of leptons and quarks with the Higgs field in the Standard Model (SM)[8, 9, 10].  [^2] When constructing the effective field theory, we will use the notation similar to the notation adopted in [11, 12] when presenting the electroweak part of the SM. This made it possible to assign masses to fermion fields, predict the strength of fermion interactions with the Higgs boson, and introduce the Cabibbo-Kobayashi-Maskawa mixing matrix. Non-Abelian gauge bosons and interactions due to the exchange of these bosons were introduced by Yang and Mills [13] when considering Heisenberg’s isotopic symmetry as a local symmetry. An attempt to interpret the emerging gauge fields, called Yang-Mills fields, as fields of vector $\rho$-mesons could not be successful due to the massless nature of the gauge fields. In the present paper, we continue research in this direction, using an analogue of the Braut-Engler-Higgs (BHE) [14, 15, 16] mechanism of spontaneous symmetry breaking in the SM to impart mass to gauge fields.
@@
One of such symmetries, which follows from QCD neglecting the current masses of quarks, is the global chiral symmetry corresponding to the $SU(2)_{L}\times SU(2)_{R}$ transformation group of left- and right-chiral quark doublets. In the Standard Model, quarks acquire masses through the BEH mechanism, quark condensate appears in complex QCD vacuum. As a result the $SU(2)_{L}\times SU(2)_{R}$ symmetry is broken, and the resulting pseudoscalar Goldstone bosons are pi-mesons with small masses of the order of 100 MeV compared to the masses of protons and neutrons. The theory of pion-nucleon and pion-meson forces was developed by Steven Weinberg [17]. In this approach the vector mesons appear as dynamical fields realising a gauge symmetry non-linearly with a number of attractive phenomenological features [18, 19, 20, 21, 22, 23, 24, 25] (see also review [26] and references therein and to it). Masses of the gauge fields are introduced by means of non-abeliean variant of the Stueckelberg mechanism [27, 28, 29].
The symmetry $SU(2)_{L}\times SU(2)_{R}$ is broken to the vector symmetry $SU(2)_{V}$. This symmetry is nothing more than isotopic symmetry, introduced at one time by Heisenberg[30]. On the basis of this symmetry, the ideas of the non-Abelian gauge invariance of Yang and Mills, and the idea of spontaneous symmetry breaking, we will construct an effective field theory of nucleon and meson interactions. The linear realization of spontaneous symmetry breaking is used similarly to the electroweak Standard Model. In the constructed effective , in particular, a prediction is obtained for the structure of diagrams that contribute to the internucleon one-boson-exchange potential(OBEP), (see, for example, CD-Bonn potential[4] and review [7]).
@@ 2 Isotopic invariance as a gauge symmetry
$\Psi_{N}(x)\rightarrow\Psi^{\prime}_{N}(x)=U(x)\Psi_{N}(x),\,\,U(x)=e^{i\alpha_{j}(x)\tau_{j}}$ (3)
@@
$\partial_{\mu}\rightarrow D_{\mu}=\partial_{\mu}-ig_{N}\rho^{j}_{\mu}\tau^{j}.$ (4)
@@
$\rho_{\mu}^{\pm}\equiv\frac{1}{\sqrt{2}}(\rho_{\mu}^{1}\mp i\rho_{\mu}^{2}),\,\,\rho_{\mu}^{0}\equiv\rho_{\mu}^{3}$ (15)
@@
$L_{\Psi}=\bar{\psi}_{p}(i\partial_{\mu}\gamma^{\mu}-M_{N}^{0})\psi_{p}+\bar{\psi}_{n}(i\partial_{\mu}\gamma^{\mu}-M_{N}^{0})\psi_{n}$ (17)
@@ 4 Electromagnetic interactions
$D_{\mu}=\partial_{\mu}-ig_{N}\rho^{j}_{\mu}\tau^{j}-ig_{Y}\frac{Y}{2}B^{Y}_{\mu},$ (27)
@@
$\rho_{\mu}^{\pm}\equiv\frac{1}{\sqrt{2}}(\rho_{\mu}^{1}\mp i\rho_{\mu}^{2}).$ (29)
@@
where the electric charges and projections of the isotopic spins of the proton and neutron are $Q_{\psi_{p}}=1$ and $Q_{\psi_{n}}=0$, $T^{3}_{\psi_{p}}=\frac{1}{2}$ and $T^{3}_{\psi_{n}}=-\frac{1}{2}$, and the constant $e$ is related to the fine structure constant $\alpha$ by the well-known equality $\alpha=\frac{e^{2}}{4\pi}$. The Gell-Mann-Nishijima relation $Q=\frac{Y}{2}+T^{3}$ for the proton $\psi_{p}$, neutron $\psi_{n}$ and $\Phi_{I}$ fields, as follows and should be from our construction, give the hypercharge value $Y=1$. Taking into account this ratio, from equation (32) we get:
@@ 5 Gauge-invariant effective interactions involving pions
$D_{\mu}^{ij}=\partial_{\mu}\delta^{ij}-ig_{N}\epsilon^{ijk}\rho_{\mu}^{k}.$ (39)
@@
$\rho_{\mu}^{\pm}\equiv\frac{1}{\sqrt{2}}(\rho_{\mu}^{1}\mp i\rho_{\mu}^{2}),\,\,\rho_{\mu}^{3}\equiv\rho_{\mu}^{0}\cos{\theta_{I}}+A_{\mu}\sin{\theta_{I}}$ (40)
$\pi^{\pm}\equiv\frac{1}{\sqrt{2}}(\pi^{1}\mp i\pi^{2}),\,\,\pi^{3}\equiv\pi^{0}$
@@ 7 Concluding remarks
The potentials can be obtained from the amplitudes given in Figs.4,5,6 by a well-known method, which is presented in detail in the appendices to review[7] for various variants of bosons exchanged between nucleons. The explicit form of the potentials following from the exchange diagrams of the presented theory is not given in this paper.

### v3

@@ 1 Introduction
One of the first theories of nucleon-nucleon interactions was the Yukawa theory [1]. Subsequently, the meson exchange potentials between nucleons began to be intensively constructed and studied. Various variants of potentials were obtained, which made it possible to understand many properties of nuclei, to quantitatively describe data on cross sections, phases, and scattering lengths. A number of specific potentials have received much attention, such as the Moscow potential [2], the Argonne potential [3] or later ones built on the basis of field theories, such as the Bonn potential [4], the Nijmegen potential [5, 6] and others. Review article [7] provides a detailed overview of the history and current state of research on internucleon potentials, in particular, the internucleon one-boson-exchange potentials (OBEP).
The idea of exchange interaction, the exchange of pi-mesons and other mesons with masses less than the mass of the nucleon, turned out to be exceptionally fruitful not only for describing nuclear forces. Yukawa-type interactions formed the basis for describing the interaction of leptons and quarks with the Higgs field in the Standard Model (SM)[8, 9, 10].  [^2] When constructing the effective field theory, we will use the notation similar to the notation adopted in [11, 12] when presenting the electroweak part of the SM. This made it possible to assign masses to fermion fields, predict the strength of fermion interactions with the Higgs boson, and introduce the Cabibbo-Kobayashi-Maskawa mixing matrix. Non-Abelian gauge bosons and interactions due to the exchange of these bosons were introduced by Yang and Mills [13] when considering Heisenberg’s isotopic symmetry as a local symmetry. An attempt to interpret the emerging gauge fields, called Yang-Mills fields, as fields of vector $\rho$-mesons could not be successful due to the massless nature of the gauge fields. In the present paper, we continue research in this direction, using an analogue of the Braut-Engler-Higgs (BEH) [14, 15, 16] mechanism of spontaneous symmetry breaking in the SM to impart mass to gauge fields.
@@
One of such symmetries, which follows from QCD neglecting the current masses of quarks, is the global chiral symmetry corresponding to the $SU(2)_{L}\times SU(2)_{R}$ transformation group of left- and right-chiral quark doublets. In the Standard Model, quarks acquire masses through the BEH mechanism, quark condensate appears in complex QCD vacuum. As a result the $SU(2)_{L}\times SU(2)_{R}$ symmetry is broken, and the resulting pseudoscalar Goldstone bosons are pi-mesons with small masses of the order of 100 MeV compared to the masses of protons and neutrons. The theory of pion-nucleon and pion-meson forces was developed by Steven Weinberg [17]. In this approach the vector mesons appear as dynamical fields realising a gauge symmetry non-linearly with a number of attractive phenomenological features [18, 19, 20, 21, 22, 23, 24, 25, 26, 27] (see also review [28] and references therein and to it). Masses of the gauge fields are introduced by means of non-abeliean variant of the Stueckelberg mechanism [29, 30, 31].
The symmetry $SU(2)_{L}\times SU(2)_{R}$ is broken to the vector symmetry $SU(2)_{V}$. This symmetry is nothing more than isotopic symmetry, introduced at one time Heisenberg[32]. On the basis of this symmetry, the ideas of the non-Abelian gauge invariance of Yang and Mills, and the idea of spontaneous symmetry breaking, we will construct an effective field theory of nucleon and meson interactions. The linear realization of spontaneous symmetry breaking is used similarly to the electroweak Standard Model. The constructed gauge invariant effective theory predicts a very specific structure of the interaction vertices of not only protons and neutrons with rho mesons, sigma mesons, pions and photons but also the self-interaction of vector gauge bosons, their interaction with scalar sigma and pseudoscale pi mesons. The Feynman rules, following from the Laragangian of the effective theory, allow to calculate from the presented Feynman diargrams single-boson exchange potentials, amplitudes, scattering cross sections, etc. using well-known procedures. These calculations and corresponding phenomenological analyses are out of the scope of the current theoretical study and are planned as the next step.
An interesting prediction of the theory is the small difference in masses between charged and neutral $\rho$ mesons, similar to the difference in masses of $W^{\pm}$ and $Z$ bosons in the Standard Model. This difference turns out to be on the order of current experimental errors, which motivates future more accurate measurements of the rho meson masses.
@@ 2 Isotopic invariance as a gauge symmetry
$\Psi_{N}(x)\rightarrow\Psi^{\prime}_{N}(x)=U(x)\Psi_{N}(x),\,\,U(x)=e^{\mathbf{i}\alpha_{j}(x)\tau_{j}}$ (3)
@@
$\partial_{\mu}\rightarrow D_{\mu}=\partial_{\mu}-\mathbf{i}g_{N}\rho^{j}_{\mu}\tau^{j}.$ (4)
@@
$\rho_{\mu}^{\pm}\equiv\frac{1}{\sqrt{2}}(\rho_{\mu}^{1}\mp\mathbf{i}\rho_{\mu}^{2}),\,\,\rho_{\mu}^{0}\equiv\rho_{\mu}^{3}$ (15)
@@
$L_{\Psi}=\bar{\psi}_{p}(\mathbf{i}\partial_{\mu}\gamma^{\mu}-M_{N}^{0})\psi_{p}+\bar{\psi}_{n}(i\partial_{\mu}\gamma^{\mu}-M_{N}^{0})\psi_{n}$ (17)
@@ 4 Electromagnetic interactions
$D_{\mu}=\partial_{\mu}-\mathbf{i}g_{N}\rho^{j}_{\mu}\tau^{j}-\mathbf{i}g_{Y}\frac{Y}{2}B^{Y}_{\mu},$ (27)
@@
$\rho_{\mu}^{\pm}\equiv\frac{1}{\sqrt{2}}(\rho_{\mu}^{1}\mp\mathbf{i}\rho_{\mu}^{2}).$ (29)
@@
where the electric charges and projections of the isotopic spins of the proton and neutron are $Q_{\psi_{p}}=1$ and $Q_{\psi_{n}}=0$, $T^{3}_{\psi_{p}}=\frac{1}{2}$ and $T^{3}_{\psi_{n}}=-\frac{1}{2}$, and the constant $e$ is related to the fine structure constant $\alpha$ by the well-known equality $\alpha=\frac{e^{2}}{4\pi}$. The Gell-Mann-Nishijima relation $Q=\frac{Y}{2}+T^{3}$ for the proton $\psi_{p}$, neutron $\psi_{n}$ and $\Phi_{I}$ fields gives the hypercharge value $Y=1$. Taking into account this ratio, from equation (32) we get:
@@ 5 Gauge-invariant effective interactions involving pions
$D_{\mu}^{ij}=\partial_{\mu}\delta^{ij}-\mathbf{i}g_{N}\epsilon^{ijk}\rho_{\mu}^{k}.$ (39)
@@
$\rho_{\mu}^{\pm}\equiv\frac{1}{\sqrt{2}}(\rho_{\mu}^{1}\mp\mathbf{i}\rho_{\mu}^{2}),\,\,\rho_{\mu}^{3}\equiv\rho_{\mu}^{0}\cos{\theta_{I}}+A_{\mu}\sin{\theta_{I}}$ (40)
$\pi^{\pm}\equiv\frac{1}{\sqrt{2}}(\pi^{1}\mp\mathbf{i}\pi^{2}),\,\,\pi^{3}\equiv\pi^{0}$
@@ 7 Concluding remarks
The potentials can be obtained from the amplitudes given in Figs.4,5,6 by a well-known method, which is presented in detail in the appendices to review[7] for various variants of bosons exchanged between nucleons. The explicit form of the potentials following from the exchange diagrams of the presented theory is not given in this paper. The derivation of symbolic expressions for the amplitudes and nucleon-nucleon potentials, the calculation of scattering cross sections and phase shifts, as well as the corresponding phenomenological analysis and comparison of the results with experimental data go beyond the scope of the presented theoretical study and are planned as next steps.